FEL Upgrade Project Weekly Brief
September 25-29, 2000


We generated some very interesting test data on the high voltage stand-off capabilities of nitrogen ion implanted stainless steel.  Field gradients of 24 MV/m were attained over a large area (100 cm2) electrode with a minimal dark current (145 pA).  These results have considerable impact on the design of new high performance electrodes for the upgrade photogun


The FEL Project received an excellent review at this week’s SURA/DOE review of Jefferson Lab’s Science and technology program.  The review panel thought that the Upgrade project was the right and proper next step after the kilowatt demo. The committee recognized our challenge to energize the user community behind scientific applications of the present and upgraded FEL. They thought the FEL had already demonstrated and had a strong potential to further enrich and diversify the  Jefferson lab mission.

Michelle Shinn represented the team at a DOD Applications of Lasers meeting which was held Sept. 25-27 at the Pennsylvania State University.  Of the 100 people attending, a good cross-section of industry and government was represented.  She distributed information about our user facility and upgrade plans; as there was a moderate amount of interest.  Michelle also spent some time with J. Cook (JTO) and J. Albertine, discussing our near-concentric optical cavity, and how to check that the coatings will meet specifications.  She also met with John Eric (Wright-Patterson AFB, who may be involved in the management of our AF funded UV project) and gave him some background on the IR Demo.

WBS 3 (Beam Physics)

   We began to organize our efforts needed to consider changes in the electrostatics of gun electrode geometry.  The upgraded gun may allow a slightly larger and different shape cathode which can help the high field performance.  We are going to start by resuscitating the original electrostatic model done by D.Engwall and work from there to evaluate potential modifications.
B.Yunn will be spearheading this task.

  Work continued on modifications of the lattice that will allow the UV chicane/wiggler extension to be added to the baseline upgrade lattice. This modification will be a component of the planned UV program which has been proposed to AFOSR for start-up funding in FY01.

WBS 4 (Injector)

The field emission test system has been running with a 9um polished N2 implanted SST cathode and a 1um polished SST anode at a 5 mm gap with excellent results.  The gap  has held off 24 MV/m with a very small dark current (145 pA).  The conditioned gap at 21MV/m is drawing 90 pA.  These results present an excellent candidate material for the upgrade gun electrode.

 We had discussions with the SRF team (J.Mammosser) and our collaborator from KEK (K.Saito) about KEK’s proposal to treat cavities by electropolishing.  Currently, we have sent one test 7-cell cavity to KEK and a second one will be sent in the next month or so.  As reported at the LINAC conference in August, KEK has produced by electropolishing several 9-cell cavities with gradients greater than 30MV/m.  KEK will make a formal proposal to Jefferson Lab to treat all eight, 7-cell ,cavities for the FEL upgrade cryomodule in exchange for our performance test data.

WBS 6 (RF)

Continued installation of RF Zones 3 & 4.  It is expected that checkout of Zone 4 will start in 1-2 weeks.  RF Control Modules will be a problem for the next 6 months or more for both zones.
We are working with Glassman HV to provide a slave unit for the Gun HVPS to achieve 10 ma beam current.

WBS 8  (I & C)

Primary focus of the week was the development and deployment of LiTaO3 pyroelectric sensors. These will be used as both power monitors and optical beam position monitors.  Five 10 mm square detectors are being prepared for power monitoring at various brewster window ports. These are position independent as well as wavelength independent for maintaining a fixed average power to a given experiment whether the beam is pulsed (to 10 microseconds) or CW. The electronics have been somewhat challenging to provide sufficient gain and response time with limiting the noise, rise times on order of a few microseconds is desired. These outputs run to the sample & hold card then to the A/D. The Optical Beam Position Monitors (OBPM) use the same type crystal configured as four quadrants of 1mm X 1mm each. The new 32 channel VME sample & hold card layout is nearly complete, the initial boards will arrive late next week and will be operational ~10/10/00 just prior to the start of the run. Progress continues to be made on providing pico-motor control for two turning mirrors in each of the six optical labs. A chassis which multiplexes one pico-motor channel to 32 individual outputs was completed this week. The VME control card is the FEL standard general purpose I/O, many thanks to Al Grippo for this software and for the OBPM code!

WBS 9 (Transport)

o Details of most of the GX Prototype parts are ready for checking.   A glitch in the TOSCA software for the 3D model of the GX dipole was discovered.  The results had an anomalous spike.  We found conductor blocks made of "bricks" were the culprit.  The conductor for the GX model was changed to a combination of arcs and straights and the model ran to completion much more rapidly with no anomalous spikes.

 Poles:  The Survey & Alignment Group performed a detailed CMM inspection of the First Article quadrant.  The inspection showed that most of the milled surfaces were within tolerance but that there was a small systematic offset from pole to back leg and most of the holes were out of tolerance.  Dave McCay and Jeff Karn are working the problem with the vendor.
 Budgetary Quote: After a meeting held with Everson last Friday, they anticipated having a completed quote by September 29th.   Measurement Probe: Deflection calculations have been made by R. Hicks using a variety of materials.
o We made two layouts integrating the wide-angle beam dump and its required shielding into the machine so Dave Douglas could continue the design iteration.
Additional responses to the business and technical questions about the
proposals for the engineering services contract were received.

WBS 10 (Wiggler)

 The Dispersion Section package is out for bid.

WBS 11 (Optics)

This week, our upgrade efforts concentrated on the following areas:
Optical diagnostics check outs mentioned in last week's report were successful, and we learned that our beam propagation data stream had undocumented characters that prevented our plotting the data.  This has been corrected and we'll try again Friday afternoon.  A few other tests on the pyroelectric detectors that will be used on the optical beam position monitoring (O-BPM) system will be conducted today.
Detailing check prints for the mirror test stand continued this week.


 We had some hardware check-out operations this week in preparation for the October User run.